sin25°cos115°=1/2(sin40°-1)

Grade:11

4 Answers

noogler
8 years ago
apply the formula 2cosAsinB=sin(A+B)-sin(A-B)
1/2(2cos115osin25o)=1/2(sin240o-sin90o)                              
 
plz check ur questin again.i think it should be  sin25°cos115°=1/2(sin240°-1)
 
approve if it is r8
Lab Bhattacharjee
8 years ago
2sin25°cos115°=sin140°-sin90°=sin(180°-40°)-1
Use sin(180°-y)=sin y



 
noogler
8 years ago
sorry for the wrong ans
 
apply the formula 2cosAsinB=sin(A+B)-sin(A-B)
1/2(2cos115osin25o)=1/2(sin140o-sin90o)
=1/2(sin(180-40)-1)=1/2(sin40o-1)                             
 
 
approve if it is r8
Ragini
3 years ago
Sin 25°. Cos 115°
Applying formula : sin A.cos B = 1/2 [ sin ( A+B) + Sin (A – B ) ]
1/2 [ sin ( 25° + 115°) + sin ( 25° – 115° ) ]
1/2 [ sin 140° – sin 90° ]
1/2 ( sin 40° – 1)